A pizza place offers 4 different cheeses and 10 different toppings. In how many ways can a pizza be
made with 1 cheese and 4 toppings?

Answer by ewatrrr(10682)

 
Hi,
A pizza place offers 4 different cheeses and 10 different toppings
In how many ways can a pizza be made with 1 cheese and 4 toppings?
4C1*10C4 = 4*210 = 820 ways